# Evaluating LLMs on microservice-based applications: how complex is your specification?

### v1

@@ 6 Related Work
Extending LLM code generation to tackle more realistic real-world problems for software engineers is an active area of research. One trend in this direction is software engineering agents, who facilitate developers in their everyday tasks. More specifically, as related to our work, is research into extending LLM code synthesis to tackle more realistic and complex scenarios. Much of this research is focussed on automatically correcting issues or extending the functionality of a solution in a GitHub repository [10, 13, 3, 30, 15]. While the fix or code extension may be relatively minor, the difficulty is in providing the LLM with the correct context, which can be non-trivial given the size of the repository. In contrast, our work focusses on taking a specification of an application that has multiple microservices, and generating code from scratch to realize the specification. The challenge in this problem is the complexity of the specification. It requires generating code that incorporates business logic, invokes the appropriate internal and external dependencies, incorporates non-functional considerations such as logging and authentication, and handles a variety of inputs and multiple error conditions.
@@
We found that for moderately difficult specifications GPT-o3-mini did quite well. However, for more difficult specifications it performed quite poorly. By examining the errors, we found specific issues that it had trouble resolving. For instance, one problem is that LLMs will sometimes hallucinate when importing packages [24, 12] . This displays itself when the LLM imports a package that does not really exist, or invokes non-existent APIs of in a package. Locating dependencies and implementing the APIs correctly is another recognized problem [14, 16, 22, 25].
@@
While other research has investigated the use of LLMs together with REST APIs, it has been primarily focused on how to generate test cases for a OpenAPI specifications [26, 6]. In contrast, our goal is to generate code for applications that express their interfaces via REST APIs. Surprisingly, we do not know of any other works that address this problem.
@@
In this paper we propose a complexity metric for MSBA specs, including dependency difficulty. Along somewhat similar lines, the authors in [32] define $6$ runnable levels to classify the scope required for a function to run successfully; e.g., from stand-alone functions to those that require knowledge of code in other source files. They also propose a metric Acc@k to measure the percent of target functions that are referenced in at least one of the $k$ LLM-generated sample codes. A similar measure, Recall@k, is defined in [34]. Our complexity metric is focussed on the specification complexity and includes additional components, such as length of the specification.
@@ 7 Conclusion
The goal of this paper is to evaluate how well LLMs do in generating code for microservice-based applications. We defined a standard template for specifying such applications, a framework for testing them, and proposed a metric for scoring the complexity of a specification; that is, the difficulty of generating code that satisfies the specification. We ran experiments on 8 microservices from 2 applications. Our results indicate, unsurprisingly, that the greater the complexity of the specification, the more difficulty the LLM has in passing unit tests for the microservices. While stronger LLMs such as GPT-3o-mini do fairly well in generating code for specifications of medium difficulty, they do very poorly on complex specifications that have more intricate business logic, more dependencies, and require authentication and database integration. We analyzed the key problems LLMs face in generating code for these specifications thereby suggesting future research directions to improve the capabilities of LLMs to generate microservice-based applications. It is yet to be seen whether stronger LLMs will require additional techniques to generate code for MSBAs or will be able to do so on their own.
@@ Appendix A Prompts
This prompt provides the MSBA description and the code generation guidelines. The MSBA descriptions for the public library and the restaurant nutrition MSBAs are too long to include here. They can be found online (see footnote 2). The code generation guidelines are given in Figure 3. The reflection prompt used to regenerate code when there are errors is presented in Figure 4.
Code Generation Guidelines- Each microservice should be written in Python.- The Python packages that need to be installed   must be listed in a comment  at the beginning of the code.- The code must be executable. Besides comments in the code, no other text  in the response.- Each microservice should store its persistent data in a Mongo database.  Assume that Mongo is available at ’mongodb://mongo:27017/’.- If a microservice requires data that another microservice is responsible for,  it invokes a GET request on that microservice to retrieve  that data when it  needs it. It does not  replicate and store the data locally.- Assume that each microservice will run in its own Docker container.- For all services, each REST request should return a status code in addition  to the JSON payload if there is one.  For successful requests, the appropriate  2XX code is returned:  201 (successful POST), 204 (successful DELETE) and  200 (success on other requests).  For unsuccessful requests, the appropriate 4XX or 5XX code is returned: 400  (Bad request), 404 (Not Found), 405 (Method Not Allowed), 415 (Unsupported  media type), and 500 (Internal Server Error).
@@
The prompt for regenerating a program after reflection is:
"When the Python Interpreter tried to execute this program, there wereerrors.  Here is the summary and instructions on how to fix these errors:{reflection}.   Re-generate the code and make sure to fix the errors.Check that you have fixed each error."
Appendix B LLM-as-a-Judge
@@
Measure the complexity of implementing the following micro-servicesapplication specification on a scale 1-5 (1 is easy, 5 is hard):<spec>The ratings and comments on the complexity of the MSBAs given in this paper are provided in Figures 5 and 6. We used Claude Sonnet 4 to provide these ratings. We note that we separately asked the same LLM to score each individual service. It was consistent for the restaurant app where it rated Authentication and Ratings at $2$ (out of $5$), the profile service at $3$ and the dishes service at $4$. However for the library app it rated Logs at $1$, Cardholders at $3$, Books at $4$, and Borrows at $5$.
Looking at this public library microservices application specification, Iwould rate the complexity as **3 out of 5** (moderate complexity).Here’s my analysis:**Factors that keep complexity moderate (not higher):**- **Clear, well-defined requirements** - The specification is detailed and    unambiguous- **Standard CRUD operations** - Most services follow basic Create, Read,    Update, Delete patterns- **Common technology stack** - Python, MongoDB, Docker, REST APIs are     well-established- **Simple data models** - Straightforward JSON documents with basic field     types- **Limited business logic** - Most complexity is in basic calculations    (fines, return dates)**Factors that add complexity:**- **Inter-service communication** - Services need to call each other    (Cardholders → Borrows, Borrows → Logs)- **External API integration** - Google Books API integration for    publisher/date lookup- **Business rule enforcement** - Preventing borrowing when 2+ books are    overdue- **Date calculations** - Fine calculations based on overdue days, return    date computation- **Error handling** - Proper HTTP status codes and exception handling across services- **Distributed data consistency** - Ensuring data integrity across multiple services    and databases**Why not higher complexity:**- No complex algorithms or data processing- No authentication/authorization requirements- No real-time or high-performance requirements- No complex integrations beyond one external API- No sophisticated business workflows**Why not lower complexity:**- Multiple interconnected services require coordination- Business logic beyond simple CRUD operations- External dependencies and error scenarios to handleThis falls squarely in the "moderate" category - more complex than a simple single-service CRUD app, but less complex than enterprise-scale distributed systems withcomplex business logic, security requirements, or performance constraints.
Figure 5: Claude Sonnet 4 comments on the complexity of the library MSBA.
@@
Appendix D The effects of temperature

### v2

@@ 7 Related Work
Extending LLM code generation to tackle realistic real-world problems for software engineers is an active area of research. One trend in this direction is software engineering agents, who facilitate developers in their everyday tasks. More directly related to our work is research in extending LLM code synthesis to tackle more realistic and complex scenarios. Much of this research is focussed on automatically correcting issues or extending the functionality of a solution in a GitHub repository [13, 17, 2, 34, 19]. While the code repair or extension may be relatively minor, the difficulty is in locating the relevant context, which can be non-trivial given the size of the repository. In contrast, our work focusses on taking a specification of an application that has multiple microservices and generating code from scratch to realize the specification. The challenge here is the complexity of the specification. It requires generating code that incorporates business logic, invokes the appropriate internal and external dependencies, incorporates non-functional considerations such as logging and authentication, and handles a variety of inputs and multiple error conditions.
@@
We found that for moderately difficult specifications GPT-o3-mini did quite well. However, for difficult specifications it performed poorly. By examining the errors we found specific issues that it had trouble resolving. Some of these problems have been identified in previous work. For instance, one issue is hallucinations in importing packages [28, 14] . This displays itself when the LLM imports a package that does not really exist, or invokes non-existent APIs of in a package. Locating dependencies and implementing the APIs correctly is another recognized problem [18, 20, 26, 29, 38]. In [10] they state that “For RWPB (real world projects) … the proportion of runtime bugs, especially incorrect arguments and incorrect boundary condition checks, is significantly higher compared to existing benchmarks.” We additionally identified type inference and datatype conversion as a problem. This issue is investigated in [22].
@@
While other research has investigated the use of LLMs together with REST APIs, it has been primarily focused on how to generate test cases for OpenAPI specifications [30, 8, 4]. In contrast, our goal is to generate code for applications that express their interfaces via REST APIs. We are unaware of other works that explicitly address this problem.
@@
In this paper we propose a complexity metric for specifications, including dependency difficulty. Along somewhat similar lines, the authors in [36] define $6$ runnable levels to classify the scope required for a function to run successfully; e.g., from stand-alone functions to those that require knowledge of code in other source files. They also propose a metric Acc@k to measure the percent of target functions that are referenced in at least one of the $k$ LLM-generated sample codes. A similar measure, Recall@k, is defined in [38]. Our complexity metric includes additional components, such as length of the specification and resource complexity. In [10] the authors conclude that “As the complexity of the benchmarks increases, the accuracy of LLMs tends to decline, especially when the lines of code and the number of APIs increase.” They also find that “Despite the large context windows of current LLMs, their capacity to handle lengthy and complex problems is still limited. They struggle to retain all problem details, and the extended code generation results in more syntactical errors like unmatched parentheses.”
@@ 8 Conclusion
The goal of this paper is to evaluate how well LLM Agents do in generating code for microservice-based applications. We defined a standard template for specifying such applications, a framework for testing them, and a metric for scoring the complexity of a specification. We ran experiments on 8 microservices from 2 applications. Our results indicate, unsurprisingly, that the greater the complexity of the specification, the greater difficulty the LLM has in passing unit tests for the microservices. While stronger LLMs such as GPT-3o-mini do fairly well in generating code for specifications of medium difficulty, they do poorly on complex specifications that have more intricate business logic, more dependencies, and require authentication and database integration. We analyzed the key problems LLM Agents face in generating code for these specifications thereby suggesting future research directions to improve the capabilities of LLM Agents to generate microservice-based applications. We showed that a fine-grained approach to code-generation, generating code per request type as opposed to code for the entire service, produces more accurate code.
@@ Appendix A Prompts
This prompt provides the MSBA description and the code generation guidelines. The MSBA descriptions for the public library and the restaurant nutrition MSBAs can be found online (see footnote 2). The code generation guidelines are given in Figure 3. The reflection and program regeneration prompts are presented in Figures 4 and 5. The initial MSBA description, the previous code generated, and the error and test messages are also supplied with these prompts.
Code Generation Guidelines- Each microservice should be written in Python.- The Python packages that need to be installed must be listed in a comment at the beginning of the code.- The code must be executable. Besides comments in the code, no other text in the response.- Each microservice should store its persistent data in a Mongo database. Assume that Mongo is available at ‘mongodb://mongo:27017/’.- If a microservice requires data that another microservice is responsible for, it invokes a GET request on that microservice to retrieve that data when it needs it. It does not replicate and store the data locally.- Assume that each microservice will run in its own Docker container.- For all services, each REST request should return a status code in addition to the JSON payload if there is one. For successful requests, the appropriate 2XX code is returned: 201 (successful POST), 204 (successful DELETE) and 200 (success on other requests). For unsuccessful requests, the appropriate 4XX or 5XX code is returned: 400 (Bad request), 404 (Not Found), 405 (Method Not Allowed), 415 (Unsupported media type), and 500 (Internal Server Error).
@@
When the Python Interpreter tried to execute this program, there wereerrors. Here is the summary and instructions on how to fix these errors:{reflection}. Re-generate the code and make sure to fix the errors.Check that you have fixed each error.
Figure 5: The code regeneration prompt
Appendix B LLM-as-a-Judge Prompt
@@
Measure the complexity of implementing the following micro-servicesapplication specification on a scale 1-5 (1 is easy, 5 is hard):<spec>We used Claude Sonnet 4 to provide these ratings. The ratings and comments provided by the LLM Judge for the restaurant MSBA is given in Figure 6. The comments for the library app can be found online.
@@
Appendix D Prompt for fine-grained microservice generation
The prompt to generate code only for a POST request for the dishes microservice is given in Figure 8. Prompts for the other request types and other services are similar.
Below is (I) a description of the microservices for a restaurant nutrition applicationand (II) code generation guidelines. Generate code for the POST request for thedishes microservice following the guidelines. When generating the code for thePOST request, you should specify the following before the generated functiondefinition:@app.route(’/dishes’, methods=[‘POST’])@authenticate(required_privilege=True)By specifying @authenticate, you need not explicitly authenticate the user.You can assume that the generated code will be embedded in a file containing thefollowing import statements:from flask import Flask, request, jsonifyfrom pymongo import MongoClientimport requestsimport jsonimport osfrom functools import wrapsYou can also assume that the generated code will be embedded in a file containingthe code:"if __name__ == ‘__main__’:"so do not generate another statement of this form.I. <MSBA spec>II. <code generation guidelines>
Figure 8: The fine-grained prompt
Appendix E The effects of temperature